OpenStudy (anonymous):

help.......... A man has two more nickels than dimes. He has $1.15 in all. How many coins of each kind does he have? Which equation can be used to solve the problem? 5x + 10(x + 2) = 115 5(x + 2) + 10x = 115 x + x + 2 = 115

OpenStudy (austinl):

Well, nickles appear to be the x value in this problem. So, it is x+how many more he has than dimes. Which one appears to satisfy that?

OpenStudy (anonymous):

is it the first one......5x+10(x+2)=115

OpenStudy (austinl):

Ok... now that I look at this.... 5 is the value for nickles 10 is the value for dimes 5x+10x=115 That means that they have equal values for each coin to add up to $1.15 But there are two more dimes than there are nickles, so, 5(x+2)+10x=115

OpenStudy (anonymous):

ok.. thankx